OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

jimthompson5910 (jim_thompson5910):

double everything in equation 1 to go from 4x + y = 8 to 8x + 2y = 16

jimthompson5910 (jim_thompson5910):

so you now have this system 8x + 2y = 16 x − 2y = 2 add the two equations to get 8x + 2y = 16 x − 2y = 2 ------------ 9x + 0y = 18 9x = 18 I'm sure you know what to do from here.

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

yep, you got it

OpenStudy (anonymous):

ty

jimthompson5910 (jim_thompson5910):

np